OpenStudy (gabebae):

Pyramid A is a square pyramid with a base side length of 14 inches and a height of 6 inches. Pyramid B has a volume of 3,136 cubic inches. How many times bigger is the volume of pyramid B than pyramid A? Give your answer as a percent. Provide an explanation and proof for your answer to receive full credit.

OpenStudy (gabebae):

This is what I have so far The volume of the first pyramid is going to be a^2 h/3 14^2 6/3 = 196*2 = 392

OpenStudy (gabebae):

@AlexandervonHumboldt2